OpenStudy (anonymous):

if the length of 1 leg of a right triangle is 5 and the hypotenuse has length 45, what id the length of the third side

OpenStudy (anonymous):

5^2 + x^2 = 45^2 25 + x^2 = 2025 x^2 = 2000 x= \[20\sqrt{5}\]

OpenStudy (anonymous):

a^2 + b^2 = c^2 is standard formula for right triangle solve for unknown leg; say b a = SQRT(c^2-b^2) = SQRT(45^2-5^2) a = 44.721359549995793928183473374626

OpenStudy (anonymous):

Pythagorean Theorem \[a^2+b^2=c^2\] a and b are the legs of a right triangle and c is the hypotenuse. So plug in what you have. let us say a=5 and c=45, therefore we need to solve for b So we have: \[25+b^2=2025\] subtract 25 from both sides \[b^2=2000\] then take the square root of both sides \[b \approx 44.7214\]